1)Evaluate the logarithms using the change of base formula. Round your result to three decimal places.
log20 5
2)Use properties of logarithms to expand the expression as a sum differences and/constant mutiple of logarithms.
log5 x^2/y^2z^3
3) Condense the expression to logarithm of a single quantity.
3log3 x+4log3 y-4log3z

Answer by stanbon(75887)

1)Evaluate the logarithms using the change of base formula. Round your result to three decimal places.
log20 5 = ln(5)/ln(20) = 0.5372
-----------------------------------------
2)Use properties of logarithms to expand the expression as a sum differences and/constant mutiple of logarithms.
log5 x^2/y^2z^3
= log5(x^2) - log5(y^2)-log5(z^3)
= 2log5(x) - 2log5(y)- 3log5(z)
------------------------------------------------
3) Condense the expression to logarithm of a single quantity.
3log3 x+4log3 y-4log3z
log3(x^3) + log3(y^4) - log3(z^4)
---
log3[x^3*y^4/z^4]
===========================================
Cheers,
Stan H.